OpenStudy (anonymous):

solve for x. sqrt2x-3 =5

OpenStudy (anonymous):



OpenStudy (anonymous):

apply the opposite so opposite of sqrt is square. 2x - 3 = 5^2 2x = 5^2 + 3 x = (5^2 + 3)/2 x = you can do this.

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

by squaring both sides you cancel out the square root on the left side.

OpenStudy (anonymous):

which just leaves 2x - 3

OpenStudy (anonymous):

25+3

OpenStudy (anonymous):

x = (5^2 + 3)/2 =(25+3)/2=28/2

OpenStudy (anonymous):

got it????????

OpenStudy (anonymous):

28/2=14

OpenStudy (anonymous):

yes that's right.

OpenStudy (anonymous):

ok..is that all?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do i solve 28/2???

OpenStudy (anonymous):

yes, so x=14.

OpenStudy (anonymous):

solve for x is like saying what is the value of x

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no problem.